# Linguistic Neuron Overlap Patterns to Facilitate Cross-lingual Transfer on Low-resource Languages

### v1

@@ 5.2 Ablation Study
In this part, we conduct ablation study to evaluate the impact of neuron probe data and the proposed HSIC similarity metric on bridge selection, using the BLI task for example.
Table 3: The impact of neuron probe data. ‘w/o *’ denotes replacing our constructed probe data with bilingual tokens extracted from FLORES+ dataset.
@@
Table 3 presents the results of ablation experiments to compare “w/o *” with our constructed probe data, where “w/o *” denotes replacing the constructed probe data with the bilingual tokens extracted from FLORES+ NLLB Team et al. (2024). Appendix C illustrates the detailed construction for “w/o *”. These results highlight impact of probing data in neuron manipulation. Table 4 in Appdedix C compares the performance of HSIC and Cosine similarity and shows the proposed HSIC helps to capture the dependency between language-overlapping neurons and specific neurons.
@@ 5.3 Overlapping Neuron Distribution
This section analyzes the distribution of overlap neurons. Figure 5 compares neurons in language pairs within the same family (e.g., Ar-He) and across families (e.g., Ar-Sw). Obviously, Ar-He shares more overlap neurons. Similar observations can be found in comparing language pairs from different source languages (Zh, Ar, Id, and En ) to a same target language He (Figure 10 in B.4) From the perspective of language families, Figure 6 examines the activated behaviors of neurons in low-resource languages. Obviously, low-resource languages within the Uralic family have the highest activation frequency, while Indo-European languages have the lowest. We hypothesize LLMs activate neurons more frequently for processing low-resource languages due to their perceived difficulty.
@@ 6 Conclusion
In this work, we explore whether sharing neurons can improve LLMs’ cross-lingual performance on low-resource languages. We propose a simple yet effective language-bridge approach with the help of neuron interpretation. To ensure accurately and fully activate overlap neurons across languages, we construct neurons probe data from the ground-truth MUSE dictionaries. By quantifying neuron similarity, we seek the optimal bridge for X-ICL and conduct extensive experiments to validate the efficacy and generalization of our approach.
@@
This work focuses on sharing neurons across languages and relies on evaluated data to validate the effectiveness of our approach. Due to the limitation of lacking evaluation benchmarks for low-resource languages, the experiments are conducted on 15 language pairs and have not been extensively validated on a wide range of low-resource languages. We select 4 typologically diverse low-resource languages from distinct families and test their performance on 2 cross-lingual tasks. Second, our study reveals that high-quality probe data is essential for accurately analyzing neurons’ behaviors of low-resource languages. The proposed linguistic distance measurement is probe-data-induced, providing qualitative insights, but lacks quantitative precision.
Acknowledgments
This work was supported by the National Social Science Foundation (No.24CYY107) and the Fundamental Research Funds for the Central Universities (No.2024TD001).
@@ Appendix A Appendix: Linguistic Similarity Based on Glottolog Phylogenetic Trees
We leverage Glottolog verion 5.1 Hammarström et al. (2023) as a foundational phylogenetic framework to calculate the linguistic similarity of human languages. It has two key steps: data preprocessing and similarity calculation.
Data Preprocessing. The preprocessing pipeline consists of three steps: 1) Glottocode Identifiers with regex pattern matching to ensure unambiguous language node identification. 2) Standardize node naming with underscores (e.g., [sini1245]→ _sini1245_), ensuring consistent formatting in downstream phylogenetic analyses. 3) Mitigate encoding conflicts through temporary file caching. These steps preserve both accurate phylogenetic tree parsing and computational compatibility.
Similarity Calculation. The proposed metric integrates two well-established phylogenetic principles from historical linguistics, which includes node distance normalization and depth-adjusted compensation.
First, building upon Wichmann & Holman’s framework for typological stability assessment Wichmann and Holman (2009), we compute the inter-language distance $d(L_{1},L_{2})$ between languages $L_{1}$ and $L_{2}$ using ETE3’s optimized tree traversal algorithms Huerta-Cepas et al. (2016) and then normalize it to make distances comparable across language families, calculated as:
@@
Appendix B Appendix: Neuron Patterns
B.1 Deactivation Overlap Neurons
Figure 7 presents overlap neurons distributions and their deactivation effects on Chinese-Hebrew BLI task.
Figure 7: Overlap neurons distributions and their deactivation effects on Chinese-Hebrew BLI task.
B.2 Linguistic Spectrum in Mistral
Figure 8: Mistral’s linguistic spectrum across 15 languages from 7 families. The color intensity represents the degree of overlap between language pairs.
B.3 Parameter $K$ Discussion
Here we discuss which $k$ middle layers should be selected to quantify linguistic similarity. According to observations in section 3.3.3, neurons in middle-layers should be prioritized over final-layers when measuring language similarity. We use embedding semantic similarity metric to determine $K$. We compute the layer-wise embedding semantic similarity between Arabic-Hebrew and Chinese-Hebrew pairs when predicting the same Hebrew token. As presented in Figure 9, we find embedding similarity is stable in the middle layers 10-21 and is not sensitive to variations in the predicted tokens. Therefore, $K$ layers is set to be 10-21 in LLaMA 3 and 15-23 in Mistral.
Figure 9: Embedding semantic similarity between Arabic-to-Hebrew and Chinese-to-Hebrew translations when predicting the same token at each layer.
B.4 Overlap Neuron Distribution
@@
Appendix C Appendix: Ablation Results
This section presents the detailed experimental ablation to evaluate the impact of neuron probe data construction and the HSIC similarity metric on bridge selection. Table 3 and Table 4 present the results of ablation experiments on the BLI task. “w/o *” denotes replacing our probe data with a simplified version based on FLORES+. For example, “w/o *” probe data in Indonesian-Hebrew is illustrated in Figure 11.
@@ Appendix C Appendix: Ablation Results
Figure 11: Example of “w/o *” probe data in Indonesian-Tagalog.
Appendix D Appendix: Neuron Semantic Analysis
In this section, we analyze the semantic similarity of overlapping neurons in two language groups: Hebrew-Tagalog-Swahili (He-Tl-Sw, different language family) and Portuguese-Spanish-Italian (pt-Es-It, same language families). For each group, we select overlapping neurons and the same number of randomly sampled neurons for comparison. We input $m$ parallel sentences and record the neuron activation frequency for each sentence, obtaining three $m\times 100$ activation matrices for overlap and random neurons. These matrices are mapped to a 2D semantic space using UMAP McInnes et al. (2018), with each point representing a neuron activated by a sentence.
As shown in Figure 12, randomly sampled neurons align with linguistic relationships. For example, random neurons of He and Tl are close, indicating semantic proximity in the same family. The overlap neurons cluster together, both within and across language families, proving that our approach can effectively capture semantic similarity.
@@ Appendix D Appendix: Neuron Semantic Analysis
Appendix E Appendix: Prompt Templates

### v2

@@ 5.3 Ablation Study
In this part, we conduct ablation study to evaluate the impact of the constructed neuron probe data and the proposed HSIC similarity metric on bridge selection, using the BLI task as an example.
Table 4: The impact of neuron probe data. ‘w/o *’ denotes replacing our constructed probe data with bilingual tokens extracted from the FLORES+ dataset.
@@
Table 4 presents the ablation results by comparing “w/o *” with our constructed probe data, where “w/o *” denotes replacing the constructed probe data with the bilingual tokens extracted from FLORES+ NLLB Team et al. (2024). The detailed construction of “w/o *” is provided in Appendix D. The results highlight the crucial impact of probe data on effective neuron manipulation. Furthermore, Table 6 in Appendix D compares HSIC with Cosine similarity, showing that HSIC better captures the dependency between language-overlapping neurons and specific neurons.
@@ 5.4 Overlapping Neuron Distribution
This section analyzes the distribution of overlap neurons. Figure 5 compares neurons in language pairs within the same family (e.g., Ar-He) and across families (e.g., Ar-Sw). Obviously, Ar-He shares more overlapping neurons. Similar observations can be found in comparing language pairs from different source languages to a same target language (Figure 10 in Appendix C.4)
From the perspective of language families, Figure 6 examines the activated behaviors of neurons in low-resource languages. Obviously, low-resource languages within the Uralic family have the highest activation frequency, while Indo-European languages have the lowest. We hypothesize LLMs activate neurons more frequently for processing low-resource languages due to their perceived difficulty.
@@ 6 Conclusion
In this work, we explore whether sharing neurons can improve LLMs’ cross-lingual performance on low-resource languages. We propose a simple yet effective language-bridge approach with the help of neuron interpretation. To ensure accurate and full activation of overlap neurons across languages, we construct probing data from the ground-truth MUSE dictionaries. By quantifying neuron similarity, we seek the optimal bridge for X-ICL and conduct extensive experiments to validate its efficacy and generalization.
@@
This work focuses on sharing neurons across languages and relies on the evaluated datasets to validate the effectiveness of our approach. Due to the lack of comprehensive benchmarks for low-resource languages, our experiments cover only $15$ language pairs and select $4$ low-resource languages from distinct families as target languages and test their performance on $4$ cross-lingual tasks. Second, our study reveals that high-quality probe data is crucial to accurately analyze neuron behaviors of low-resource languages, while the proposed linguistic distance measurement is probe-data-induced, offering qualitative but not quantitative insights. Finally, although bridge selection should ideally follow linguistic phylogeny, we aim to select bridges that LLMs can best exploit, inevitably reflecting their training biases.
@@ Appendix A Appendix: Linguistic Similarity Based on Glottolog Phylogenetic Trees
We leverage Glottolog version 5.1 Hammarström et al. (2023) as a foundational phylogenetic framework to calculate the linguistic similarity of human languages. It has two key steps: data preprocessing and similarity calculation.
Data Preprocessing. The preprocessing pipeline consists of three key steps: 1) Locate Glottocode identifiers with regex pattern matching to ensure unambiguous language node identification. 2) Standardize node naming with underscores (e.g., [sini1245] $\rightarrow$ _slaini1245_), ensuring consistent formatting in downstream phylogenetic analyses. 3) Mitigate encoding conflicts through temporary file caching. These steps preserve accurate parsing of phylogenetic tree while maintaining computational compatibility.
Similarity Calculation. The proposed metric integrates two well-established principles from historical linguistics: node distance normalization and depth-adjusted compensation.
First, building upon Wichmann & Holman’s framework for typological stability assessment Wichmann and Holman (2009), we compute the inter-language distance $d(L_{1},L_{2})$ between languages $L_{1}$ and $L_{2}$ using ETE3’s optimized tree traversal algorithms Huerta-Cepas et al. (2016). We then normalize the distance to make it comparable across language families, calculated as:
@@
Appendix B Appendix: Supplementary Cross-lingual Tasks
To further verify robustness, we conduct additional experiments using the bridge languages selected by our method on downstream tasks, including Cross-Lingual Question Answering (CLQA) and Cross-Lingual Natural Language Inference (XNLI), as shown in Table 5. Due to the limited availability of cross-lingual benchmarks covering our target low-resource languages (e.g., Tagalog), the evaluation is restricted to $6$ language pairs for CLQA and $3$ pairs for XNLI.
Table 5: Evaluation on CLQA and XNLI cross-lingual tasks. Bold highlights improved performance.
CLQA XNLI
Zero-shot Ours Zero-shot Ours
Zh-Ja 42.80 -0.60 - -
Zh-Sw 38.20 +5.20 35.90 +1.30
Ar-Ja 41.60 +6.20 - -
Ar-Sw 34.20 +2.40 35.50 +3.70
Id-Ja 34.60 +3.60 - -
Id-Sw 34.80 +5.20 36.60 +0.60
Appendix C Appendix: Neuron Patterns
C.1 Deactivation Overlap Neurons
Figure 7 presents the distribution of overlap neurons and their deactivation effects on the Chinese-Hebrew (Zh-He) BLI task.
Figure 7: Overlap neuron distributions and their deactivation effects on the Chinese-Hebrew BLI task.
C.2 Linguistic Spectrum in Mistral
Figure 8: Mistral’s linguistic spectrum across $15$ languages from 77 families. The color intensity represents the degree of overlap between language pairs.
C.3 Parameter $K$ Discussion
Here we discuss which $k$ middle layers should be selected to quantify linguistic similarity. According to observations in section 3.3.3, neurons in middle layers should be prioritized over final layers when measuring language similarity. We use the embedding semantic similarity metric to determine $K$. For example, we analyze Arabic-Hebrew and Chinese-Hebrew translation pairs by prompting LLMs with the same semantical words in Arabic and Chinese to generate the corresponding Hebrew translation. We then compute the layer-wise embedding semantic similarity between the two pairs and identify layers in which this similarity is insensitive to variations in the predicted tokens, reflecting the inherent distance between the languages. As shown in Figure 9, we find embedding similarity is stable in the middle layers 10-21 of LLaMA 3 and is not affected by token-level variations. Therefore, $K$ layers is set to be 10-21 in LLaMA 3 and 15-23 in Mistral.
Figure 9: Embedding semantic similarity between Arabic-Hebrew and Chinese-Hebrew translations when predicting the same token at each layer of LLaMA 3.
C.4 Overlap Neuron Distribution
@@
Appendix D Appendix: Ablation Results
This section presents the detailed experimental ablation to evaluate the impact of neuron probe data construction and the HSIC similarity metric on the bridge language selection. Table 4 and Table 6 present the results of ablation experiments on the BLI task. “w/o *” denotes replacing our probe data with a simplified version based on FLORES+. For example, Figure 11 illustrates the construction of “w/o *” probe data for Indonesian-Hebrew.
@@ Appendix D Appendix: Ablation Results
Figure 11: Example of “w/o *” probe data in Indonesian-Hebrew.
Appendix E Appendix: Neuron Semantic Analysis
In this section, we analyze the semantic similarity of overlapping neurons across two language groups: Hebrew-Tagalog-Swahili (He-Tl-Sw, from different language families) and Portuguese-Spanish-Italian (Pt-Es-It, from the same language family). Our goal is to examine whether overlapping neurons cluster together within the same language family or also across different families.
For comparison, we select $100$ overlapping neurons and randomly sample the same number of non-overlapping neurons for comparison. We feed the model with $m$ parallel sentences and then record the neuron activation frequency for each sentence, obtaining three $m\times 100$ activation matrices for both overlapping and random neurons. These matrices are mapped to a 2D semantic space using UMAP McInnes et al. (2018), with each point representing a neuron activated by $m$ sentences from a language. Colored circles of red, yellow, and blue denote overlapping neurons and colored $\times$ symbols of red, yellow, and blue denote random neurons.
As presented in Figure 12, the overlapping neurons, whether identified within the same family or across language families, cluster closely together. This suggests that our approach can effectively identify neurons that encode genuine shared semantics across languages. In addition, random neurons tend to align with linguistic relationships, clustering when feeding LLMs tokens from topologically related languages, such as Portuguese (Pt), Spanish (Es), and Italian (It), while remaining dispersed when feeding LLMs tokens from distant languages, like Hebrew (He), Tagalog (Tl), and Swahili (Sw).
(a) Neuron semantic similarity in Pt-Es-It
(b) Neuron semantic similarity in He-Tl-Sw
@@ Appendix E Appendix: Neuron Semantic Analysis
Appendix F Appendix: Prompt Templates